There are three consecutive integers such that the sum of the first and second integers decreased by the third integer will result in 68. What are the integers?
Answer by checkley71(8403)

LET X, X+1 & X+2 BE THE 3 NUMBERS.
X+(X+1)-(X+2)=68
X+X+1-X-2=68
X-1=68
X=68+1
X=69 ANSWER FOR THE FIRST NUMBER
X+1=70
X+2=71
PROOF
69+70-71=68
139-71=68
68=68